OpenStudy (anonymous):

Which of the following would result in ? 3x^2 A.3xy divide xy b. 7 x^y - 4x^2y c. xy+2xy D.18x^3y^2/6xy^2

OpenStudy (sh3lsh):

A would result in 3 because the x and the y would cancel out. B stays the same because the exponents on the variables are different (x is different than x^2) C Would result in 3xy (you just add the numbers if the exponents are the same) D Would result in 3x^2 (the 18/6 = 3 ) (x^3/x = x^2) and (y^2/y^2 cancels out)

OpenStudy (anonymous):

so b ?

OpenStudy (sh3lsh):

D

OpenStudy (anonymous):

oh i got it thanks